Ringmeup, Inc., had net income of $183,600 for the year ended December 31, 2016. At the beginning of the year, 39,000 shares of common stock were outstanding. On May 1, an additional 17,000 shares were issued. On December 1, the company purchased 4,400 shares of its own common stock and held them as treasury stock until the end of the year. No other changes in common shares outstanding occurred during the year. During the year, Ringmeup, Inc., paid the annual dividend on the 7,000 shares of 4.85%, $100 par value preferred stock that were outstanding the entire year. Required: Calculate basic earnings per share of common stock for the year ended December 31, 2016. (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

Earnings attributable to common stock holders

= Net income - preferred dividend

= 183,600 - (7,000 x 100 x 4.85%)

= 183,600 - 33,950

= $149,650

Weighted average number of shares outstanding

= (39,000 x 11/12) + (17,000 x 8/12) + [(39,000 - 4,400) x 1/12]

= 35,750 + 11,333.33 + 2,883.33

= 49,966.67

Basic EPS = ​​​​​​Earnings attributable to common stockholders/weighted average shares

= 149,650/49,966.67

= $2.99